Friday, August 3, 2007
MEB DUCT CLEANING
- Ceiling Installation
- Landis continues to work of the punchlist items
- 1st floor air handlers were disconnected and moved to their new
temporary locations. Re-installation will take place next week.
- Rough-in for light, fixtures and outlets continues.
- Door frame and sheetrock installation is complete.
- Carpet Installation
- Carpet Installation continues in Library
- Awaiting viable sample results
- Zones 4C and 4D are complete.
- Unit 4-B was vacuumed, cleaned, and coated. The ducts will be coated this evening.
- Ordes pulled wires from the disconnect, mounted at the entrance to the lot at
the Medical School Building, to the guard booth in the L&M lot.
- Siemens continues to remove old. fire alarm wiring and devices.
- Installed last of the garden edging for Lion's Eye Building on the S. Prieur St. side.